) Give an example to show that if the convergence of 􏰄 an is conditional and (bn) is a bounded

sequence, then 􏰄 anbn may diverge.


1
Expert's answer
2022-05-16T12:25:36-0400

ANSWER :"\sum_{n=1}^{\infty}a_{n}=\sum_ {n=1}^{\infty}\frac{(-1)^{n } }{n},\, \, b_{n}=(-1)^{n}"

EXPLANATION The series "\sum_ {n=1}^{\infty}\frac{(-1)^{n } }{n}" is an alternating series, since the sequence "a_{n}=(-1)^{n}c_{n}" and "c_{n}= \frac {1}{n

}" is decreasing , "\lim _{n\rightarrow\infty}c_{n}=0" . The series "\sum_ {n=1}^{\infty}\frac{1 }{n}" diverges, because it is a "p-" series for "p=1" , hence the series "\sum_ {n=1}^{\infty}\frac{(-1)^{n } }{n}" converges conditionally . The sequence "b_{n}=(-1)^{n}" is a bounded sequence , since "-1\leq b_{n}\leq 1" for all "n\in\N" ("b_{2n }=1,\, b_{2n-1}=-1" ). Since "a_{n}\cdot b_{n}=\frac {1}{n}" , then the series "\sum_{n=1}^{\infty}a_{n}\cdot b_{n}=\sum_ {n=1}^{\infty}\frac{1 }{n}" diverges.